OpenStudy (anonymous):

-y+10=25 Btw, Mrbonez321, this is my last one for tonight! LOL

OpenStudy (anonymous):

awww man... :( -y+10=25 -10 -10 -y=15 -- --- -1 -1 y=-15

OpenStudy (anonymous):

subtract 10 from both sides. -y=15. Multiply both sides by -1 to get a positive variable. & i agree with Mrbonez321 "awww man... :("

OpenStudy (anonymous):

-y+10=25 -10 -10 -y= 15 -- -- -1 -1 y= -15

OpenStudy (anonymous):

-y+10=25 -10 -10 -y=15 -- --- -1 -1 y=-15

OpenStudy (anonymous):

Its okay! I'll definitely have more Tuesday. And thank you all! You guys are too smart. :I